# Flow-Modulated Scoring for Semantic-Aware Knowledge Graph Completion

### v2

@@ 3.2.1. Semantic Context Learning Module
For a given triplet $(h,r,t)$ in Knowledge Graphs, its semantic context offers valuable cues for identifying valid links from $h$ to $t$ (Wang et al., 2021). Traditional node-based message passing is applicable to general graphs and has led to numerous variants, such as GCN (Kipf and Welling, 2016), GraphSAGE (Hamilton et al., 2017), and GAT (Vaswani, 2017). However, when applied to knowledge graphs, we observe that edges themselves also possess corresponding features (i.e., relation embeddings $x^{e}$), and the number of relation types is significantly smaller than the number of entities. Modeling message passing over relations is thus more efficient and provides more indicative semantic cues.
@@
To this end, we adopt an improved alternating relational message passing scheme, with its core idea inspired by PathCon (Wang et al., 2021), and enhance its expressive power by introducing a semantic-aware neighbor selection strategy. Specifically, when generating a message $m_{i}^{v}$ for a node $v$ by aggregating information from its edges $\mathcal{E}(v)$, naively aggregating information from all edges can introduce noise and may lead to information dilution or the over-smoothing problem due to the inclusion of semantically irrelevant edges. To address this challenge, we introduce a Top-K selection mechanism based on semantic similarity. Prior to aggregation, we first evaluate the semantic relevance between node $v$ (represented by its own representation $h^{v}$) and each of its neighboring edges $e$ (represented by its current state $s_{i}^{e}$). This is achieved by a learnable mapping function $g(\cdot)$ that projects $h_{v}$ and $s_{i}^{e}$ into a shared latent semantic space. In this space, we compute their similarity score using an energy-model inspired scoring function:
(3) $\text{Score}(v,e)=\exp(-\|g(h^{v})-g(s^{e}_{i})\|^{2}/\tau),$
@@
where $\|\cdot\|^{2}$ is the squared Euclidean distance and $\tau$ is a temperature hyperparameter. Edges with higher scores, deemed more semantically relevant to node $v$, are selected to form the Top-K set $\mathcal{E}^{\prime}_{K}(v)\subseteq\mathcal{E}(v)$.
@@
The states of these selected Top-K edges, $\{s^{e}_{i}\}_{e\in\mathcal{E}^{\prime}_{K}(v)}$, are first mean-aggregated to form a consolidated neighborhood representation:
(4) $\bar{s}^{e}_{K,i}(v)=\text{Mean}\left(\{s^{e}_{i}\}_{e\in\mathcal{E}^{\prime}_{K}(v)}\right).$
Figure 3. The architecture of attention aggregator.
@@
This aggregated edge representation $\bar{s}^{e}_{K,i}(v)$, along with the node’s own representation $h^{v}$, is then processed by our specialized node message aggregator. This aggregator (as shown in Figure 3), employs a multi-head attention mechanism. It transforms $h^{v}$ and $\bar{s}^{e}_{K,i}(v)$, calculates head-wise gating scores based on their interaction, applies these gates to the transformed $\bar{s}^{e}_{K,i}(v)$, and then combines this gated representation with the transformed $h^{v}$ before a final linear projection. This yields a more semantically focused node message $m_{i}^{v}$:
(5) $m^{v}_{i}=\text{AttentionAggregator}\left(h^{v},\bar{s}^{e}_{K,i}(v)\right).$
The impact of the semantic scoring function and the aggregator selection will be further discussed in subsequent ablation studies.
Next, we generate a message $m_{i}^{e}$ for each edge $e$. This message is derived from the node messages $m_{i}^{v}$ and $m_{i}^{u}$ received from its two endpoint nodes $v$ and $u$ as follows:
(6) $m^{e}_{i}=\text{MLP}([m^{v}_{i},m^{u}_{i}]).$
Finally, the newly generated edge message $m_{i}^{e}$ is used to update the hidden state of edge $e$. The update rule is as follows:
(7) $s^{e}_{i+1}=\sigma\left(\text{Linear}\left(\left[s^{e}_{i},m^{e}_{i}\right]\right)\right)\equiv\sigma\left(\left[s^{e}_{i},m^{e}_{i}\right]\cdot W_{i}+b_{i}\right),$
where $\sigma$ is a non-linear activation function, $\left[s^{e}_{i},m^{e}_{i}\right]$ represents the concatenation of the previous edge state and the current edge message, and $W^{i}$ and $b^{i}$ are the learnable weight matrix and bias term for layer $i$, respectively. The initial hidden state of an edge is defined by its intrinsic feature embedding $s^{e}_{0}=x^{e}\in\mathbb{R}^{d_{e}}$.
@@
In this manner, we not only leverage the structure and features of edges in the KG but also, through the semantic-aware neighbor selection mechanism and the subsequent attention-based aggregation, enable the model to more accurately capture and propagate context information most relevant to the specific relation prediction task.
3.2.2. Condition Flow-Matching Module
@@
Prior research typically predicts static association scores directly utilizing entity embeddings. Although our Semantic Context Learning module is able to generate more informative entity representations, it is still insufficient to fully characterize the inherent dynamic transformation properties of the relationships. To address this limitation, we have designed the Conditional Flow Matching Module.
The core of this module is learning a dynamic vector field $\boldsymbol{v}_{\theta}(t,\boldsymbol{x})$ that depends on a pseudo-time variable $t\in[0,1]$ and the current state $\boldsymbol{\mathit{x}}\in\mathbb{R}^{\scriptstyle\mathit{d}}$. This vector field characterizes the instantaneous direction and velocity along the smooth evolution path of the entity representation, conditioned on a given context $\boldsymbol{z}$, as it transitions from an initial state h (i.e., at $\mathit{t}=0$) to a target state t (i.e., at $\mathit{t}=1$). It is important to note that the variable t does not represent physical time but rather serves as a pseudo-time dimension parameterizing this transformation process.
@@
Specifically, drawing upon the principles of Conditional Flow Matching  (Tong et al., 2023b), we employ a continuous flow to model the transformation from the head entity embedding $\boldsymbol{\mathit{h}}^{*}$ to the tail entity embedding $\boldsymbol{\mathit{t}}^{*}$. (Note: The asterisk (*) is used here to distinguish entity representations from the pseudo-time variable $t$. This notation will be maintained without repeated clarification). This process can be formalized as:
[…]
@@
where $\mathit{p_{t}(\boldsymbol{\mathit{x}})}$ represents the marginal probability path, $q(\boldsymbol{\mathit{z}})$ is the sampling distribution of the semantic context $\boldsymbol{z}$, and the conditional probability path $p_{t}(\boldsymbol{\mathit{x}}|\boldsymbol{\mathit{z}})$ denotes the probability path given $\boldsymbol{z}$. Correspondingly, the marginal vector field $u_{t}(\boldsymbol{\mathit{x}})$ that drives the evolution of $p_{t}(\boldsymbol{\mathit{x}})$ can be expressed as an expectation of the conditional vector field $u_{t}(\boldsymbol{\mathit{x}}|\boldsymbol{\mathit{z}})$ with respect to the prior distribution $q(\boldsymbol{\mathit{z}})$, weighted by the posterior distribution $p(\boldsymbol{\mathit{z}}|t,\boldsymbol{\mathit{x}})$:
(9) $u_{t}(\boldsymbol{\mathit{x}})=E_{q(\boldsymbol{\mathit{z}})}\left[u_{t}(\boldsymbol{\mathit{x}}|\boldsymbol{\mathit{z}})\frac{p_{t}(\boldsymbol{\mathit{x}})}{p_{t}(\boldsymbol{\mathit{x}}|\boldsymbol{\mathit{z}})}\right],$
where the conditional vector field is defined as $u_{t}(\boldsymbol{\mathit{x}}|\boldsymbol{\mathit{z}})=\frac{d\boldsymbol{\mathit{x}}}{dt}$. Consequently, under suitable conditions, $u_{t}(\boldsymbol{\mathit{x}})$ generates the probability flow path $p_{t}(\boldsymbol{\mathit{x}})$. This formulation allows us to circumvent the need for explicit modeling of the marginal probability path $p_{t}(\boldsymbol{\mathit{x}})$.
@@
Although we aim to recover $u_{t}(\boldsymbol{\mathit{x}})$, its direct computation is generally difficult. This is primarily because the computation of $p_{t}(\boldsymbol{\mathit{x}})$ requires integration. Considering the relative tractability of the conditional vector field $p_{t}(\boldsymbol{\mathit{x}}|\boldsymbol{\mathit{z}})$, we adopt an indirect learning strategy, and the model is enabled to directly learn an approximation of the vector field $u_{t}(\boldsymbol{\mathit{x}}|\boldsymbol{\mathit{z}})$. We assume $v_{\theta}:\mbox{\small$[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$}$ is a time-dependent vector field parameterized by a neural network with parameters $\theta$. We define the following objective function:
(10) $\mathcal{L}_{cfm}(\theta)=\mathbb{E}_{t,q(\boldsymbol{\mathit{z}}),p_{t}(\boldsymbol{\mathit{x}}\mid\boldsymbol{\mathit{z}})}\left\|v_{\theta}(t,\boldsymbol{\mathit{x}})-u_{t}(\boldsymbol{\mathit{x}}|\boldsymbol{\mathit{z}})\right\|_{2}^{2}.$
@@
Based on this, the model learns $v_{\theta}(t,\boldsymbol{x})$. It can serve as a dynamic adjustment factor to modulate the base static score assigned to the corresponding entity pair. Different choices of $q(\boldsymbol{z})$ and $p_{t}(\boldsymbol{x}|\boldsymbol{z})$ determine the different focuses of the model.
In this study, we primarily employ two Conditional Flow Matching paradigms, each based on different assumptions about the underlying dynamics:
Condition Streaming Matching(CSM)
@@
This paradigm is groun-ded in Optimal Transport theory (Schrijver, 2003) and aims to model the least-cost transformation path from head entity distribution to tail entity distribution. The condition $\boldsymbol{z}$ is defined as: $\boldsymbol{z}=(\boldsymbol{h}^{*},\boldsymbol{t}^{*})$. The sampling distribution for the context $\boldsymbol{z}$, denoted as $q(\boldsymbol{z}$):
[…]
@@
where $\rho_{H}$ and $\rho_{T}$ represent the marginal distributions of the head and tail entity embeddings, respectively, and $\Pi(\rho_{H},\rho_{T})$ is the set of all joint probability distributions. This design aims to guide the model to learn a flow field aligned with the static optimal transport path. In practice, we employ the Minibatch Optimal Transport method for efficient approximation. Thus, the model tends to learn representations that can be transformed into one another with a low transmission cost. We define the conditional probability path and its corresponding conditional vector field as follows:
(12) $p_{t}(\boldsymbol{x}|\boldsymbol{z})=\mathcal{N}(\boldsymbol{x}\mid(1-t)\boldsymbol{h}^{*}+t\boldsymbol{t}^{*},\sigma^{2}\mathbf{I}),$
(13) $u_{t}(\boldsymbol{x}|\boldsymbol{z})=\boldsymbol{t}^{*}-\boldsymbol{h}^{*},$
@@
where $p_{t}(\boldsymbol{x}|\boldsymbol{z})$ is a Gaussian distribution, and the noise standard deviation $\sigma$ controls the stochasticity of the probability path. A moderate value of $\sigma$ can enhance model robustness. Through CSM, model can learn more direct transformation paths between entities.
Condition Flow Brownian Bridge Matching(CFBBM)
This matching paradigm draws upon Schrödinger Bridge theory (Schrödinger, 1932; Tong et al., 2023a) and aims to model the stochastic evolution process between $\boldsymbol{h}^{*}$ and $\boldsymbol{t}^{*}$. Its core objective is to find a path probability distribution that is closest to standard Brownian motion, subject to given their distribution constraints. $\boldsymbol{z}$ is the same as that in the previous paradigm, $q(\boldsymbol{z})$ is designed as an entropy-regularized optimal transport plan:
[…]
where $\lambda=2\sigma^{2}$ and $H(\pi)$ is the entropy of the joint distribution $\pi$. This formulation seeks to minimize the transport cost while simultaneously maximizing entropy (making the transport plan as diffuse as possible), with the trade-off controlled by the regularization parameter $\lambda$. With the introduction of entropy regularization, it makes the computation of $q(\boldsymbol{z})$ more feasible for large-scale datasets. It can alleviate the high computational complexity of the previous paradigm. Under this paradigm, the conditional probability path and its corresponding conditional vector field are defined as follows:
(15) $p_{t}(\boldsymbol{x}|\boldsymbol{z})=\mathcal{N}(\boldsymbol{x}\mid(1-t)\boldsymbol{t}^{*}+t\boldsymbol{h}^{*},t(1-t)\sigma^{2}\mathbf{I}),$
(16) $u_{t}(\boldsymbol{x}|\boldsymbol{z})=\frac{1-2t}{2t(1-t)}(\boldsymbol{x}-((1-t)\boldsymbol{t}^{*}+t\boldsymbol{h}^{*}))+(\boldsymbol{h}^{*}-\boldsymbol{t}^{*}),$
where $t(1-t)\sigma^{2}\mathbf{I}$ reflects the characteristic diffusion properties of a Brownian bridge process. Through this matching strategy, the model is guided to learn a dynamic flow field that simulates the stochastic evolutionary path connecting the embeddings of two entities. This facilitates the capture of potentially more complex and non-linear stochastic evolutionary relationships between entities.
3.2.3. Module Integration and Overall Scoring Process
In this section, we will detail how the two aforementioned modules are integrated to achieve comprehensive relation prediction.

### v3

@@ 3.2.1 Semantic Context Learning
For a given triplet $(h,r,t)$ in a Knowledge Graph (KG), its semantic context offers valuable cues for identifying valid links [19]. Traditional node-based message passing, as seen in GraphSAGE [66], GAT [67] and RED-GNN [26], is applicable to general graphs. However, when applied to KGs, we observe that edges themselves possess rich features (i.e., relation embeddings $x_{e}$), and the number of relation types is significantly smaller than the number of entities. Consequently, modeling message passing directly between relations (edges) is not only more efficient but also provides more indicative semantic cues. As illustrated in Figure 1, our reasoning about Iron Man’s ”ally” status is predicated on these relational attributes, rather than on the intrinsic properties of the entities themselves.
To this end, we propose a relational message passing scheme where edges update their states by aggregating information from neighboring edges. This approach treats the KG’s relational structure as the primary computational graph. However, naively aggregating information from all neighboring edges can introduce noise and lead to information dilution or over-smoothing, especially when some neighbors are semantically irrelevant to the target relation.
@@
To address this challenge, we introduce a Top-K selection mechanism based on semantic similarity between edges. Specifically, when updating the state of a central target edge $e_{c}$, we first evaluate its semantic relevance to each neighboring edge $e_{n}$ in its local neighborhood $\mathcal{N}(e_{c})$. This is achieved by projecting their current states, $s^{e_{c}}_{i}$ and $s^{e_{n}}_{i}$, into a shared latent semantic space using a learnable mapping function $g(\cdot)$ [68]. In this space, we compute their similarity score using an energy-model inspired scoring function:
$\text{Score}(e_{c},e_{n})=\exp(-\|g(s^{e_{c}}_{i})-g(s^{e_{n}}_{i})\|^{2}/\tau),$ (3)
@@
where $\|\cdot\|^{2}$ is the squared Euclidean distance and $\tau$ is a temperature hyperparameter. Edges with higher scores, deemed more semantically relevant to the central edge $e_{c}$, are selected to form the Top-K set $\mathcal{N}_{K}(e_{c})\subseteq\mathcal{N}(e_{c})$.
@@
The states of these selected Top-K edges, $\{s^{e_{n}}_{i}\}_{e_{n}\in\mathcal{N}_{K}(e_{c})}$, are first mean-aggregated to form a consolidated neighborhood representation:
$\bar{s}^{\mathcal{N}_{K}}_{i}(e_{c})=\text{Mean}\left(\{s^{e_{n}}_{i}\}_{e_{n}\in\mathcal{N}_{K}(e_{c})}\right).$ (4)
@@
This aggregated neighborhood representation $\bar{s}^{\mathcal{N}}_{i}(e_{c})$, along with the central edge’s own state $s^{e_{c}}_{i}$, is then processed by our specialized attention aggregator to compute the updated edge state for the next layer, $s^{e_{c}}_{i+1}$. This aggregator employs a multi-head attention mechanism that functions as a gating system. It transforms both $s^{e_{c}}_{i}$ and $\bar{s}^{\mathcal{N}}_{i}(e_{c})$, calculates head-wise gating scores based on their interaction, applies these gates to the transformed neighborhood representation, and then combines this gated information with the transformed self-representation before a final projection. This process yields a semantically focused update:
$s^{e_{c}}_{i+1}=\text{AttentionAggregator}\left(s^{e_{c}}_{i},\bar{s}^{\mathcal{N}_{K}}_{i}(e_{c})\right).$ (5)
The initial state of an edge is defined by its intrinsic feature embedding, $s^{e}_{0}=x^{e}\in\mathbb{R}^{d_{e}}$. This iterative update is performed for a predefined number of layers. This process ultimately generates a contextual representation for the central entity, which encapsulates the relational information within its multi-hop neighborhood. The impact of the semantic scoring function and aggregator choice will be further discussed in our ablation studies (subsection 4.4).
@@
In this manner, our model directly leverages the relational structure of the KG. Through the semantic-aware edge neighbor selection and the subsequent attention-based aggregation, the model accurately captures and propagates context information most relevant to the specific relation prediction task, performing reasoning directly in the relation space.Notably, this process avoids the substantial computational overhead of encoding entity embeddings by not leveraging any entity-specific features.
3.2.2 Condition Flow-Matching Module
@@
Prior research typically predicts static association scores directly utilizing entity embeddings. Although our Semantic Context Learning module is able to generate more informative entity representations, it is still insufficient to characterize the dynamics of relationships. To address this limitation, we have designed the Conditional Flow Matching Module. It constructs a conditional flow-based flow matching mechanism, incorporating the semantic context in which entities are situated as conditional variables into the relationship modeling process.
The core of this module is learning a dynamic vector field $\bm{v}_{\theta}(t,\bm{x})$ that depends on a pseudo-time variable $t\in[0,1]$ and the current state $\bm{\mathit{x}}\in\mathbb{R}^{\scriptstyle\mathit{d}}$. The reason behind our proposal of the pseudo-time variable lies is the fact that the dynamics of relationships are inherently intertwined with the temporal dimension—evolution of relationships (in the real world it is mostly accompanied by the passage of time). However, the construction of traditional knowledge graph datasets has limitations: they do not explicitly incorporate temporal annotation information. Instead, such temporal attributes are latent in the overall distribution of massive triple data, making them difficult to be directly captured and utilized. To address this, we introduce $t$ as a bridge. It does not directly replicate physical time, but rather externalizes the originally concealed temporal properties through parameterization. This treatment makes the model to more accurately capture the intrinsic logic of relationship evolution over ”time”. This learned vector field characterizes the instantaneous direction and velocity along the evolution path from the head to the tail entity embedding.
@@
Specifically, drawing upon the principles of Conditional Flow Matching [63], we employ a continuous flow to model the evolution from the head entity embedding $\bm{\mathit{h}}^{*}$ to the tail entity embedding $\bm{\mathit{t}}^{*}$ [^1] The asterisk (*) is used here to distinguish entity representations from the pseudo-time variable $t$. This notation will be maintained without repeated clarification. This process can be formalized as:
$\mathit{p}_{\mathit{t}}(\bm{\mathit{x}})=\int\mathit{p}_{\mathit{t}}(\bm{\mathit{x}}|\bm{\mathit{z}})\mathit{q}(\bm{\mathit{z}})\mathit{d}\bm{\mathit{z}},$ (6)
@@
where $\mathit{p_{t}(\bm{\mathit{x}})}$ represents the marginal probability path, $q(\bm{\mathit{z}})$ is the sampling distribution of the semantic context $\bm{z}$, and the conditional probability path $p_{t}(\bm{\mathit{x}}|\bm{\mathit{z}})$ denotes the probability path given $\bm{z}$. Correspondingly, the marginal vector field $u_{t}(\bm{\mathit{x}})$ that drives the evolution of $p_{t}(\bm{\mathit{x}})$ can be expressed as an expectation of the conditional vector field $u_{t}(\bm{\mathit{x}}|\bm{\mathit{z}})$ with respect to the prior distribution $q(\bm{\mathit{z}})$, weighted by the posterior distribution $p(\bm{\mathit{z}}|t,\bm{\mathit{x}})$:
$u_{t}(\bm{\mathit{x}})=E_{q(\bm{\mathit{z}})}\left[u_{t}(\bm{\mathit{x}}|\bm{\mathit{z}})\frac{p_{t}(\bm{\mathit{x}}|\bm{\mathit{z}})}{p_{t}(\bm{\mathit{x}})}\right],$ (7)
where $u_{t}(\bm{\mathit{x}}|\bm{\mathit{z}})=\frac{d\bm{\mathit{x}}}{dt}$. Consequently, under the initial condition $p_{t}(\bm{\mathit{x_{0}}})$, $u_{t}(\bm{\mathit{x}})$ generates the probability flow path $p_{t}(\bm{\mathit{x}})$. We formally establish this result as follows.
Proof.
To prove that the vector field $u_{t}(\bm{x})$ generates the probability path $p_{t}(\bm{x})$, we must show that they satisfy the continuity equation: $\frac{\partial p_{t}(\bm{x})}{\partial t}+\nabla_{\bm{x}}\cdot\left(u_{t}(\bm{x})p_{t}(\bm{x})\right)=0.$ We begin by taking the time derivative of the marginal probability path defined in Eq. 6:
$\frac{\partial p_{t}(\bm{x})}{\partial t}=\frac{\partial}{\partial t}\int p_{t}(\bm{x}|\bm{z})q(\bm{z})\,d\bm{z}\\ =\int\frac{\partial p_{t}(\bm{x}|\bm{z})}{\partial t}q(\bm{z})\,d\bm{z},$ (8)
where the second of these equal signs is guaranteed by the Leibniz Rule. By definition, the conditional path $p_{t}(\bm{x}|\bm{z})$ is generated by the conditional vector field $u_{t}(\bm{x}|\bm{z})$. Therefore, it satisfies its own continuity equation:
$\frac{\partial p_{t}(\bm{x}|\bm{z})}{\partial t}=-\nabla_{\bm{x}}\cdot\left(u_{t}(\bm{x}|\bm{z})p_{t}(\bm{x}|\bm{z})\right).$ (9)
Substituting this into our expression gives:
$\frac{\partial p_{t}(\bm{x})}{\partial t}$ $=\int-\nabla_{\bm{x}}\cdot\left(u_{t}(\bm{x}|\bm{z})p_{t}(\bm{x}|\bm{z})\right)q(\bm{z})\,d\bm{z}$ (10)
$=-\nabla_{\bm{x}}\cdot\int u_{t}(\bm{x}|\bm{z})p_{t}(\bm{x}|\bm{z})q(\bm{z})\,d\bm{z}.$ (11)
In the second step, we moved the divergence operator $\nabla_{\bm{x}}\cdot$ outside the integral, as the integration is with respect to $\bm{z}$. Now, let’s examine the definition of the marginal vector field $u_{t}(\bm{x})$ from Eq. 7. By multiplying both sides by $p_{t}(\bm{x})$, we get:
$u_{t}(\bm{x})p_{t}(\bm{x})=\int u_{t}(\bm{x}|\bm{z})p_{t}(\bm{x}|\bm{z})q(\bm{z})\,d\bm{z}.$ (12)
The right-hand side is precisely the term inside the integral in our derivation. Substituting this back, we obtain:
$\frac{\partial p_{t}(\bm{x})}{\partial t}=-\nabla_{\bm{x}}\cdot\left(u_{t}(\bm{x})p_{t}(\bm{x})\right).$ (13)
Rearranging the terms yields the continuity equation:
$\frac{\partial p_{t}(\bm{x})}{\partial t}+\nabla_{\bm{x}}\cdot\left(u_{t}(\bm{x})p_{t}(\bm{x})\right)=0.$ (14)
Since the marginal probability path $p_{t}(\bm{x})$ and the marginal vector field $u_{t}(\bm{x})$ satisfy the continuity equation, and $p_{t}(\bm{x})$ satisfies the initial condition $p_{0}(\bm{x})$ by construction, it follows that $u_{t}(\bm{x})$ is indeed the generator of the probability path $p_{t}(\bm{x})$. ∎
@@
Although we aim to recover $u_{t}(\bm{\mathit{x}})$, its direct computation is generally difficult. Considering the relative tractability of the conditional vector field $p_{t}(\bm{\mathit{x}}|\bm{\mathit{z}})$, we adopt an indirect learning strategy, and the model is enabled to directly learn an approximation of the vector field $u_{t}(\bm{\mathit{x}}|\bm{\mathit{z}})$. We assume $v_{\theta}:\mbox{\small$[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$}$ is a time-dependent vector field parameterized by a neural network with parameters $\theta$. We define the following objective function:
$\mathcal{L}_{cfm}(\theta)=\mathbb{E}_{t,q(\bm{\mathit{z}}),p_{t}(\bm{\mathit{x}}\mid\bm{\mathit{z}})}\left\|v_{\theta}(t,\bm{\mathit{x}})-u_{t}(\bm{\mathit{x}}|\bm{\mathit{z}})\right\|_{2}^{2}.$ (15)
@@
Based on this, the model learns $v_{\theta}(t,\bm{x})$. It can serve as a dynamic adjustment factor to modulate the static base score assigned to the pair of corresponding entities.
@@
Then, we draw on the optimal transport theory [69], and aims to model the least-cost evolution path from head entity distribution to tail entity distribution. $q(\bm{z}$) defined as:
$q(\bm{z})=\pi(\bm{z})=\underset{\pi\in\Pi(\rho_{H},\rho_{T})}{\text{argmin}}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\left\lVert\bm{x}-\bm{y}\right\rVert_{2}^{2}\,d\pi(\bm{x},\bm{y}),$ (16)
@@
where $\rho_{H}$ and $\rho_{T}$ represent the marginal distributions of the head and tail entity embeddings, respectively, and $\Pi(\rho_{H},\rho_{T})$ is the set of all joint probability distributions. This design aims to guide the model to learn a flow field aligned with the static optimal transport path. The conditional probability path and its corresponding conditional vector field are defined as in the paper [69]:
$p_{t}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x}\mid(1-t)\bm{h}^{*}+t\bm{t}^{*},\sigma^{2}\mathbf{I}),$ (17)
$u_{t}(\bm{x}|\bm{z})=\bm{t}^{*}-\bm{h}^{*},$ (18)
@@
where $p_{t}(\bm{x}|\bm{z})$ is a Gaussian distribution, and the noise standard deviation $\sigma$ controls the stochasticity of the probability path. In particular, as $\sigma\to 0$, The marginal vector field can be regarded as a generative model of $\bm{h}^{*}$. Proof is given in the Theorem. A moderate value of $\sigma$ can enhance model robustness. Through CSM, model can learn more direct evolution paths between entities.
Theorem 1.
The marginal $p_{t}$ corresponding to $q(z)=q(\bm{h}^{*})q(\bm{t}^{*})$ and the $p_{t}(x|z)$, $u_{t}(x|z)$ in (17) and (18) has boundary conditions $p_{1}=q_{1}\ast\mathcal{N}(x\mid 0,\sigma^{2})$ and $p_{0}=q_{0}\ast\mathcal{N}(x\mid 0,\sigma^{2})$, where $\ast$ denotes the convolution operator.
Proof.
We start with (6) to show the result. We note that $q(z)=q(\bm{h}^{*},\bm{t}^{*}))=q(\bm{h}^{*})q(\bm{t}^{*})$, so :
$p_{t}(x)$ $=\int p_{t}(x\mid z)q(z)\,dz$
$=\int\mathcal{N}\bigl{(}x\mid(1-t)\bm{h}^{*}+t\bm{t}^{*},\sigma^{2}\bm{I}\bigr{)}q\bigl{(}(\bm{h}^{*},\bm{t}^{*})\bigr{)}d(\bm{h}^{*},\bm{t}^{*})$
$=\iint\mathcal{N}\bigl{(}x\mid(1-t)\bm{h}^{*}+t\bm{t}^{*},\sigma^{2}\bm{I}\bigr{)}q(\bm{h}^{*})q(\bm{t}^{*})\,d\bm{h}^{*}d\bm{t}^{*}$
evaluated at $i=0,1$ respectively. Therefore, at $t=0$,
$p_{0}(x)$ $=\iint\mathcal{N}\bigl{(}x\mid\bm{h}^{*},\sigma^{2}\bigr{)}q(\bm{h}^{*})q(\bm{t}^{*})\,d\bm{h}^{*}d\bm{t}^{*}$
$=\int\mathcal{N}\bigl{(}x\mid\bm{h}^{*},\sigma^{2}\bigr{)}q(\bm{h}^{*})\,d\bm{h}^{*}$
$=q(\bm{h}^{*})*\mathcal{N}\bigl{(}x\mid 0,\sigma^{2}\bigr{)}.$
This is also true for $t=1$. ∎
3.2.3 Overall Scoring Process and Joint Training
Having detailed the Semantic Context Learning module and the Conditional Flow-Matching module for modeling dynamic evolution, this section will now describe how they are integrated into a unified scoring and training framework. The process unfolds in two sequential steps to produce a final, context-sensitive, and dynamically-aware score.